A certain particle has a weight of 20 N at a point where the acceleration due to 
gravity is 9.8 m/s2
. (a) What are the weight and mass of the particle at a point where the 
acceleration due to gravity is 4.9 m/s2
? (b) What are the weight and mass of the particle if it 
is moved to a point in space where the gravitational force is zero?

Grade:12th pass

3 Answers

Bishu Raj
4 years ago
given 
W = 20N 
We know W =ma
               20=m9.8
                m=  100/49 
a) now if g=4.9,then
               W=ma 
                W=(100/49)*4.9
                 W=10N
 
b)  if g=0 then Weight of particle is also 0
    but mass will not change
           
Vikas TU
4 years ago
 
Dear student 
The explanation for second part :
The weight of an object is a force. It is the force with which a body is attracted toward Earth or another celestial body. This means that when you are in space, away from Earth, objects do not weight anything since they do not feel gravitational attraction to the Earth.
What objects have though in space is mass. This is because mass is defined as the amount of material an object contains, and that doesn't change whether the object is on Earth, on the Moon, or anywhere in space
Advik
3 years ago
W=m*g
20N=m*10
20/10=m
m=2kg
a) now g=5m\s
    W=m*g
    W=2*5
    W=10N
    And mass remains tha same =2kg  
B) when g =0
     Mass remains same=2kg
     W=m*g
     W=2*0
     Weight=0